Which is the iterative rule for the sequence: 15.5, 13, 10.5, 8, 5.5, 3, …?
(Points : 1)
a. an = 15.5 − 2.5n
b. an = 18 – 2.5n
c. an = 15.5 + 2.5n
d. an = 18 + 2.5n

Answer by stanbon(75887)

Which is the iterative rule for the sequence: 15.5, 13, 10.5, 8, 5.5, 3, …?
(Points : 1)
a. an = 15.5 − 2.5n
b. an = 18 – 2.5n
c. an = 15.5 + 2.5n
d. an = 18 + 2.5n
----
You know the 1st term has n = 1
Looking at the 1st term (15.5) you can see that "a" would be less than
15.5 for n = 1. So "a" is not the answer.
---
"c" cannot be the answer because when n = 1, 15.5+2.5n is greater than 15.5.
----
"d" is greater than 18 for n = 1 so "d" is not the answer.
---
Ans: The answer is "b"
Check it out.
============
Cheers,
Stan H.
---------------